how do you figure these out? thank you.
Compute tangent of the angle A= arctan(7/1)+arctan(5/4). Round to .001
Find the Re (w)where w is the positive solution to y2 =6+3i
Answer #1

Given A = arctan (7) + arc tan) We know arc tan (1) = 81.87 arċ tanca) - 51:34 so A. = 81.874 51.34 A = 133.21 both sides byGiven ²=6+3i let z = atbi 4-6 +2abi = 6+3 Catbi)²= = 6+3i a +(bij +2 (a) Cbi) = 6731 [:12-1] comparing like terms on both sid